# Driving by Hybrid Navigation: An Online HD-SD Map Association Framework and Benchmark for Autonomous Vehicles

### v2

@@ Appendix D Metric Details
In the main article, we present a narrative explanation of the Association P-R accompanied by a schematic diagram. To elucidate the calculation of Association P-R more thoroughly, we include the pseudo-code for computing Association P-R, as depicted in Alg. 1.
Algorithm 1 Evaluate Association P-R
@@
15:                  Update TP/FP based on accuracy vs $acc$
@@
[…]
$\begin{split}&\text{A-P}^{50:95}=\sum_{th\in T}\text{A-P}^{th},\quad\text{A-R}^{50:95}=\sum_{th\in T}\text{A-R}^{th}\\ &\text{A-F1}^{th}=\frac{2\text{A-P}^{th}\cdot\text{A-R}^{th}}{\text{A-P}^{th}+\text{A-R}^{th}},\quad\text{A-F1}^{50:95}=\frac{2\text{A-P}^{50:95}\cdot\text{A-R}^{50:95}}{\text{A-P}^{50:95}+\text{A-R}^{50:95}}\end{split}$ (8)
where $th$ are the thresholds in association P-R as $[0.5:0.05:0.95]$ (10 thresholds).
Appendix E Model Details
@@
E.1 Path-aware attention
@@
E.2 Spatial attention
@@
E.3 Post Process
In the main manuscript, we offer a narrative explanation of the post-processing. To enhance clarity, we also present a mathematical formulation of the post-processing details. Let $\mathcal{R}$ denote the road as the vocabulary in the traditional beem search with size $|\mathcal{R}|$, and $k$ represent the beam width. At each step $t$, the algorithm maintains a set $B_{t}$ of candidates path $k$, each associated with a score $s(h)$ defined as the sum of logarithmic conditional probabilities. The search begins by selecting the token $w^{*}$ with the maximum initial probability $P(w|x)$ given as input $x$, forming the singleton initial set:
@@ E.3 Post Process
At iteration $t\geq 1$, each sequence $h\in B_{t-1}$ generates $2|\mathcal{R}|$ candidates by appending a token $w\in\mathcal{R}$ to either the left ($w\cdot h$) or right ($h\cdot w$) of $h$, forming the expanded candidate set:
@@
The process terminates at a predefined maximum length $T$ or when all sequences emit an end-of-sequence token, with the final output $\hat{h}$ selected as:
@@
Appendix F Implement Details
F.1 Model Settings
@@
Tab. 8 summarizes the architectural configurations of our proposed MAT variants (MAT-T, MAT-S, MAT-M, MAT-L). All variants adopt identical channel dimensions ([96,192,384,768,1536]), attention head counts ([4,4,8,8,8]), spatial curve orders (["z", "z-trans", "hilbert", "hilbert-trans"]) and hybrid attention mechanisms combining spatial attention (SA) and path-aware attention (PA). In particular, parameters such as patch sizes (1024), MLP ratios (matching spatial curve orders), and stochastic depth rates (0.3) are uniformly inherited across architectures, reflecting ablation study results that optimized these values for balanced accuracy-latency trade-offs. A distinctive design choice lies in the shuffling strategy, where MAT-T/S/M/L progressively refine the shuffle order to enhance token mixing in spatial attention, aligning with their increasing computational budgets. This structured configuration hierarchy enables systematic evaluation of model capacity versus efficiency, as validated by the ascending La / ms metrics (34 $\rightarrow$ 70) corresponding to deeper transformer layers.
Parameter MAT-T MAT-S MAT-M MAT-L
Blocks [2, 2, 2, 2, 2] [4, 4, 4, 4, 4] [4, 4, 4, 8, 4] [4, 4, 4, 12, 4]
@@
Latency/ms 34 56 63 70
Table 8: Model settings.
@@
learning rate 1e-5 epochs 50
block lr scaler 1e-5 warmup epochs 2
criteria CrossEntropy, CTC Loss
@@
Table 9: Train Configuration and Data augmentations.
F.2 Train Configuration
The details of the implementation are summarized in Tab. 9. Our training protocol employs the AdamW optimizer with a base learning rate of $2\times 10^{-3}$ and a cosine learning rate decay, operating in mini-batches of size 12. The weight decay regularization is set to $5\times 10^{-3}$, while the block-wise learning rate scaling ($10^{-1}$ factor) is applied to stabilize the propagation of the gradient across the transformer layers. The training process spans 50 epochs with a 2-epoch warm-up phase for learning rate initialization. Model optimization combines CrossEntropy loss for classification tasks and CTC loss for sequence alignment objectives.
F.3 Data Augmentations
For data augmentation as shown in Tab. 9, we implement a series of randomized transformations that include axis-aligned rotation around the z-axis with ±$\pm 1^{\circ}$ angular variation at a 50% application probability, isotropic scaling within the range $[0.9,1.1]$, random flipping with equal probability 50%, point cloud jittering characterized by $\sigma=0.005$ and a clip limit of 0.02, and grid sampling with spatial discretization parameters set to $[0.1,0.1,\pi/16]$. These enhancement strategies were systematically validated through ablation studies to optimize the balance between model accuracy and computational efficiency while ensuring robustness to input variations.

### v3

@@ E.1 Details of NR-PR
In the main article, we present a narrative explanation of the Navigation Refinement P-R accompanied by a schematic diagram. To elucidate the calculation of Navigation Refinement P-R more thoroughly, we include the pseudo-code for computing Navigation Refinement P-R, as depicted in Alg. 1.
Algorithm 1 Evaluate Navigation Refinement P-R
@@
15:      Update TP/FP based on accuracy vs accacc
@@
Furthermore, the formula for $\text{NR-P}^{50:95},\text{NR-R}^{50:95}$ and $\text{NR-F1}^{th},\text{NR-P}^{50:95}$is as follows:
$\begin{split}&\text{NR-P}^{50:95}=\sum_{th\in T}\text{NR-P}^{th},\quad\text{NR-R}^{50:95}=\sum_{th\in T}\text{NR-R}^{th}\\ &\text{NR-F1}^{th}=\frac{2\text{NR-P}^{th}\cdot\text{NR-R}^{th}}{\text{NR-P}^{th}+\text{NR-R}^{th}},\quad\text{NR-F1}^{50:95}=\frac{2\text{NR-P}^{50:95}\cdot\text{NR-R}^{50:95}}{\text{NR-P}^{50:95}+\text{NR-R}^{50:95}}\end{split}$ (7)
where thth are the thresholds in association P-R as $[0.5:0.05:0.95]$ (10 thresholds). Additionally, in the validation set, because NR-R is always 1, we use the NR-P value directly as NR-F1.
E.2 Curvature-based Weighted Metric
Although the length-interval stratification in the proposed NR-PR metric effectively mitigates bias towards short paths, it does not explicitly account for the geometric complexity of roads. In real-world datasets, straight roads are statistically more frequent than complex curved roads, potentially masking model deficiencies in handling high-curvature topologies. To address this and evaluate the model’s fairness across diverse road types, we introduce a Curvature-based Weighted Metric.
Complexity Measure. Defining complexity via average angular changes can be highly sensitive to point sampling rates. Therefore, we employ the Discrete Fréchet Distance between the actual path and the straight line connecting its endpoints as a robust measure of geometric complexity. A higher Fréchet distance indicates a greater deviation from a straight line, representing higher curvature or irregularity.
Figure 8: Statistics of Fréchet distances across all paths in the OMA dataset. (1) Histogram of Fréchet distances. A substantial number of paths exhibit Fréchet distances concentrated within the range of $[0,1]$, revealing a pronounced long-tail distribution in the overall data. (2) Cumulative probability of Fréchet distances. Consistent with (1), the distribution displays a significant long-tail characteristic, underscoring the importance of block-wise statistics for the NR-PR metric.
Stratification Strategy. As shown in Fig. 8, similar to length-based approach, we adopt a stratified statistical method to handle the long-tailed distribution of road complexity. We identify the 95th percentile of complexity scores in the dataset as an upper bound and uniformly divide the range $0,\text{95th percentile}]$ into 10 bins. The final metric is computed by averaging the NR-F1 scores across these bins, ensuring that complex road geometries contribute equally to the final score, rather than being overshadowed by the dominant straight roads.
Quantitative Analysis. We evaluated the MAT model on the OMA dataset using this weighted metric. Table 9 illustrates the impact of stratification on the overall score. As the number of bins increases—forcing the metric to weigh complex roads equally to straight ones—the overall performance drops significantly (e.g., from 81.6% to 60.3% on the Val set). This confirms that the dataset is dominated by simple geometries where the model performs exceptionally well, masking the challenges posed by complex topologies in unweighted metrics.
Table 9: Impact of Complexity Stratification on NR-F1. The metric score decreases as we enforce equal weighting for complex road types (increasing NN), revealing the dominance of simple roads in the unweighted score.
   Complexity Bins (NN)    1    2    5    10
   Val Set    81.6    70.2    62.7    60.3
   Test Set    51.8    38.7    33.5    32.5
Table 10 details the performance in specific complexity intervals. The results reveal a strong negative correlation between road complexity and association precision. On simple straight roads (Bin 1), the model achieves high precision (87.2% on Val). However, performance degrades drastically on high-complexity paths (dropping to 43.1% in Bin 10). This trend is even more pronounced in the Test set, indicating that while MAT is robust, geometric complexity remains a significant challenge for map association tasks. This analysis serves as a complement to the length-based metric, offering a more comprehensive view of the robustness of the model.
Table 10: Performance per Complexity Bin ($N=10$). Bin 1 represents straight roads, while Bin 10 represents high-curvature/irregular roads. The results indicate a performance degradation as geometric complexity increases.
Bin Index 1 2 3 4 5 6 7 8 9 10
Val Set 87.2 63.2 56.7 50.7 49.9 49.1 46.5 45.2 42.7 43.1
Test Set 57.9 50.4 32.0 21.2 16.8 14.7 10.7 10.6 10.4 11.6
Appendix F Method
@@
F.1 Path-aware attention
@@
It should be highlighted that the model organizes paths internally to prevent overly lengthy paths. If a path surpasses the patch size, it is divided into several groups, where each group (aside from the final one) matches the patch size. PA’s worst-case time complexity is $O(kn^{3})$, where $k$ is the patch size and $n$ is the number of tokens, if and only if $\frac{2}{3}$ centerline or road are either starting points or ending points. Clearly, this is highly improbable in real-world scenarios, and its average time complexity remains $O(k^{2}n)$, which is linear.
F.2 Spatial attention
@@
During the group stage, tokens are partitioned based on the sorted sequence, each partition matching the patch size. Following this, the model executes self-attention computations within these partitions. Similar to PA, the self-attention (SA) algorithm has a time complexity of $O(kn)$, where $k$ is the patch size and $n$ is the number of token, indicating that the complexity of SA remains linear.
F.3 Post Process
In the main manuscript, we offer a narrative explanation of the post-processing. To enhance clarity, we also present a mathematical formulation of the post-processing details. Let $\mathcal{R}$ denote the road as the vocabulary in the traditional beam search with size $|\mathcal{R}|$, and let $k=4$ represent the width of the beam. At each step $t$, the algorithm maintains a set $B_{t}$ of candidates path $k$, each associated with a score $s(h)$ defined as the sum of logarithmic conditional probabilities. The search begins by selecting the token $w^{*}$ with the maximum initial probability $P(w|x)$ given as input $x$, forming the singleton initial set:
@@ F.3 Post Process
In iteration $t\geq 1$, each sequence $h\in B_{t-1}$ generates $2|\mathcal{R}|$ candidates by appending a token $w\in\mathcal{R}$ to the left ($w\cdot h$) or right ($h\cdot w$) of $h$, forming the expanded candidate set:
@@
The process ends at a predefined maximum length $T$ or when all sequences emit an end-of-sequence token, with the final output $\hat{h}$ selected as:
@@
Appendix G Ablation Study
For the hyperparameters we proposed, additional ablation studies were executed to exhibit their stability.
Table 11: Ablation study of patch size of PA and SA.
Patch Size 64 256 1024 2048 $\infty$
NR-F150:95 77.8 78.4 78.7 78.5 78.4
Latency/ms 77 68 70 72 72
Table 12: Ablation study of group method of PA.
Group Method N.G Category Path
NR-F150:95 78.0 78.2 78.7
Latency/ms 75 73 70
Table 13: Ablation study of Input size of SD Map
    Input size of SD Map     NR-F150:95     Latency/ms
    $60\times 60$m     78.6     67
    $90\times 90$m     78.6     69
    $120\times 120$m     78.7     69
    $150\times 150$m     78.7     70
Table 14: Ablation study of $\delta$
$\delta$ NR-F150:95 Latency
1m 74.6 65
0.5m 78.1 68
0.1m 78.7 70
0.01m 78.7 79
Table 15: Ablation study of $\beta_{ce}$ and $\beta_{ctc}$ in loss function
$\beta_{ce}:\beta_{ctc}$ NR-F150:95 Latency
1:1 68.4 70
1:0.1 78.5 70
1:0.01 78.7 70
1:0.001 78.4 70
1:0 78.4 70
Table 16: Ablation study of space curve in SA.
Spatial Curve NR-F150:95 Latency/ms
Z 77.6 70
Z + TZ 78.1 70
H + TH 78.2 70
Z + TZ + H + TH 78.7 70
Table 17: Ablation study of beam width $k$ in post process.
Beam Width NR-F150:95 Latency/ms
1 78.2 69
2 78.6 70
4 78.7 70
8 78.7 78
G.1 Hyper Parameter
Patch Size of PA and SA. Tab. 12 shows that precision plateaus at patch size 256 ($\text{NR-F1}^{50:95}=78.4\%$) but increases slightly at 1024 ($0.3\%$) with a latency trade-off (+2+2 ms).
Group Method of PA. For PA grouping (Tab. 12), path-based grouping surpasses non-grouping ($0.7\%$) and category-based baselines ($0.5\%$), likely due to reduced cross-path interference.
Input size of SD map. Table 13 presents the results of the adjustment of the SD map input size. The findings reveal that alterations in the SD input size exert minimal influence on NR-F1 ($\leq 0.1$).
Sampling size of SA. Table 15 shows the findings for the spatial sample size $\delta$. When $\delta$ is set at 1m, there is a drop in model accuracy, which may be attributed to the challenge of differentiating nearby lane lines with such a substantial sampling size. When $\delta$ is below 0.5m, the accuracy of the model is decreased (<0.6<0.6). Decreasing the interval further to 0.01 does not improve the accuracy of the model.
Hyperparameter of loss function. As illustrated in Table 15, the ablation studies for $\beta_{ce}$ and $\beta_{ctc}$ indicate that high beta values make CTC the main loss, which reduces the precision of the model. In contrast, lower beta values allow CE loss to predominate, leading to stabilized NR-F1 (<0.3<0.3).
Space Curve of SA. Table 17 displays the result of ablation studies focusing on spatial curves in SA. The findings suggest that adding more diverse spatial curves improves the model’s performance. It is crucial to emphasize that including multiple spatial curve types does not affect model latency, given that each layer utilizes only a single spatial curve. Thus, the variety of spatial curve types does not increase the frequency of spatial curve sorting.
Beam width of post-process. Table 17 shows the results of the ablation test regarding the hyperparameter of beam width $k$ during post-processing. The study reveals that setting $k$ to 4 results in the best trade-off between accuracy and processing time.
Length intervals of NR P-R. To further investigate the influence of the path length distribution on the NR-F1 metric, we performed a stratified evaluation based on length intervals. The motivation for introducing length intervals stems from our observation that within the dataset, short-distance paths occur far more frequently than long-distance ones, and their corresponding F1 scores tend to be higher. Without appropriate handling, this imbalance could bias the overall NR-F1 score toward performance on short-distance paths. Inspired by the size‑based categorization strategy adopted in the MS COCO benchmark lin2014microsoft, we categorize paths into discrete length intervals, compute the F1 score within each interval, and subsequently aggregate the results into a unified NR-F1 measure. This procedure ensures that each length range contributes proportionally to the final metric, thus mitigating the dominance of short‑distance paths in the evaluation. As shown in Tab. 18, omitting this stratification (that is, computing the metric on all paths without length‑based grouping) produces a substantially inflated overall score. In contrast, increasing the number of length intervals causes the aggregated metric to gradually decrease and eventually converge to a stable value. These findings suggest that the length interval-based computation allows NR-F1 to provide a more balanced and fair evaluation on both the short- and long-distance paths, resulting in a more comprehensive evaluation of the model’s ability to capture relational associations.
Distance Threshold of NR P-R. Tab. 19 highlights the effect of various distance thresholds on the final metrics in NR P-R Step 1. These thresholds originate from those used in Reachability P-R. It is important to note that Val uses the ground-truth OP map for predictions, thus all roads are deemed true roads, while the distance threshold applies solely to the test set. Ablation experiments reveal that distance thresholds significantly influence the accuracy outcome of NR-F1. With a strict threshold of 0.5m, most of the lane lines on the OP map are regarded false detections, as they do not meet the requirement. In contrast, a more lenient threshold of 2.0 m allows most lane lines to be perceived as correct, bringing the NR-F1 to a value comparable to that of the Val set. This also indicates that there is no domain difference between the Test and Val sets; discrepancies in results are due to the threshold settings. However, considering that the accuracy of the lane regression is also pivotal, we opted for a balanced threshold of 1.0m to judge whether the OP map correctly represents a lane.
Table 18: Ablation study of length intervals of NR P-R.
Number of length intervals 1 2 3 6 10 15
NR-F150:95 / Val 81.8 79.0 78.5 78.2 78.7 78.7
NR-F150:95 / Test 51.6 47.6 47.2 45.3 44.9 45.0
Table 19: Ablation study of distance threshold of NR P-R.
   Distance threshold    0.5m    1.0m    1.5m    2.0m
   NR-F150:95 / Val    78.7    78.7    78.7    78.7
   NR-F150:95 / Test    6.6    45.0    68.2    72.4
Table 20: Ablation study of robustness to imperfect SD Maps on OMA Val / Test set with varying noise intensities. The metric reported is NR-F150:95.
Noise Type Low Medium High
Global Shift (10%, 20%, 50%) 77.5 / 43.3 76.0 / 42.9 74.6 / 41.4
Element Noise (5%, 10%, 20%) 78.3 / 44.8 77.5 / 44.3 77.5 / 44.1
Element Absence (10%, 12%, 30%) 77.5 / 44.0 76.4 / 43.7 75.5 / 43.1
G.2 Robustness to Imperfect SD Maps
In real-world applications, Standard Definition (SD) maps often suffer from imperfections caused by GPS localization errors, sensor noise, or map construction delays. To quantitatively evaluate the robustness of MAT and the proposed metrics against such imperfect inputs, we conducted additional ablation studies on the OMA Validation set.
We categorized SD map imperfections into three distinct types and injected varying levels of noise (Low, Medium, High) into the input SD maps. The noise generation protocols are defined as follows:
  • Global Shift: Simulates GPS localization errors by applying a uniform distribution offset relative to the global map range. We tested noise ratios of 10%, 20%, and 50%. Crucially, for each SD map sample, the same random offset vector is applied to all road vectors, preserving the internal shape but shifting the absolute position.
  • Element Noise: Simulates geometric construction errors by applying random perturbations relative to the global map range. We tested noise ratios of 5%, 10%, and 20%. Unlike Global Shift, the noise here is generated independently for each vector within the SD map, resulting in geometric jitter and shape distortion.
  • Element Absence: Simulates map construction incompleteness by randomly masking out road vectors. We evaluated omission ratios of 10%, 20%, and 30%, where vectors are randomly discarded from the SD map input.
The results, measured by the NR-F150:95 score, are summarized in Tab. 20. The results demonstrate the resilience of the proposed method under varying noise conditions:
  • Robustness to Shift: Even under a high global shift of 50% , the model maintains reasonable performance (74.6 / 41.4). This indicates that our Path-Aware Attention (PA) successfully captures the relative topological structure of the road network, reducing dependency on absolute coordinate alignment.
  • Robustness to Element Noise: The performance remains highly stable, dropping only marginally from 78.3 / 44.8 to 77.5 / 44.1 despite a 20% geometric jitter. This suggests that the Spatial Attention (SA) mechanism effectively aggregates global contextual information, thereby mitigating the impact of local geometric inaccuracies.
  • Robustness to Missing Elements: The model exhibits strong adaptability to missing data, maintaining a score of 75.5 / 43.1 even when 30% of the SD map vectors are absent. The Transformer architecture effectively infers associations based on the remaining context and the overall graph topology, ensuring that the association process does not fail catastrophically when individual segments are missing.
These experiments confirm that MAT maintains strong robustness when dealing with imperfect SD maps, effectively handling geometric misalignment, shape distortion, and topological incompleteness inherent in real-world navigation tasks.
G.3 Robustness under Environmental Noise and Occlusion
To evaluate the robustness of MAT against sensor failures caused by environmental factors and severe occlusions, we conducted extended experiments focusing on two aspects: quantitative evaluation on corrupted inputs and qualitative analysis of occlusion scenarios.
Robustness against Environmental Noise. We utilized nuScenes-C xie2025benchmarking, a benchmark designed to assess robustness against input corruptions, to simulate sensor degradation. Specifically, we used MapTRv2 maptrv2 to perform inference on the nuScenes-C validation set with pon split, generating OP maps under three specific weather conditions: Rain, Fog and Dark (Night). These conditions were categorized into three difficulty levels: Easy, medium, and Hard followed by the nuScenes-C setting.
Crucially, we adopted a zero-shot inference setting: the MAT model used for association was trained solely on clean data and was not retrained on these corrupted samples. This setting serves as a rigorous stress test for the generalizability of the model.
Table 21 reports the quantitative results with NR-F150:95 score. As observed, MAT maintains robust performance across all conditions. Even in “hard” settings, where visual features are significantly compromised, the performance drop is minimal. For example, in Fog scenarios, the metric only decreases from 44.2 (Easy) to 42.1 (Hard). This stability demonstrates that, by taking advantage of the topological priors of SD maps, MAT effectively mitigates the impact of sensor noise.
Table 21: Zero-shot robustness evaluation on nuScenes-C under different environmental corruptions.
    Condition     Easy     Mid     Hard
    Rain     44.7     42.1     40.6
    Fog     44.2     43.4     42.1
    Dark     44.0     43.8     41.6
Figure 10: Visualization of OP map degradation samples from nuScenes-C. The columns represent different environmental conditions (Rain, Fog, Dark), and the rows represent increasing difficulty levels (Easy, Mid, Hard). These noisy inputs simulate realistic sensor failures.
Figure 11: Qualitative results under “Hard” environmental conditions. The figure compares the ground truth (GT) with MAT’s predictions under severe Rain, Fog, and Dark conditions. The model demonstrates strong zero-shot robustness, effectively associating noisy OP lanes with SD paths.
Figure 12: Qualitative results under “Hard” environmental conditions and severe occlusion. Similar to the setup in Fig. 10, this figure displays the inference results under Rain, Fog, and Dark scenarios, but specifically highlights a case of severe vehicle occlusion. As observed, the model maintains robust lane reconstruction capabilities even when the visual information is corrupted by extreme weather or significantly obstructed by other vehicles.
Visualization of Environmental Degradation. To intuitively display the challenge, Fig. 10 illustrates the synthesized degradation samples of the OP maps at different difficulty levels. Fig. 11 and Fig. 12 further compare the inference results under the “Hard” setting. Despite the severe noise in the input OP maps (e.g. missing boundaries or phantom lanes due to fog/rain), MAT successfully recovers the correct lane topology by aligning it with the SD map.
Analysis of Severe Occlusion. While the explicit quantitative evaluation of occlusion is constrained by the lack of ground-truth annotations, we analyze it physically. Vehicle occlusion typically manifests itself in OP maps as lane break or missing segments. Since MAT is explicitly designed to model global topology, it is inherently robust to such defects.
Fig. 12 presents a case study of severe vehicle occlusion. As shown in the visualization, MapTRv2 manages to reconstruct a coherent global road structure despite significant visual blockage by vehicles. We attribute this capability to the temporal modeling design of the upstream perception model (MapTRv2), where single-frame occlusion does not disrupt the map modeling derived from the overall temporal sequence. Taking advantage of the inherent robustness of MapTRv2 against occlusion, our downstream MAT model is consequently largely unaffected by dynamic vehicle occlusion, ensuring stable association performance.
Appendix H Visualization
H.1 Attention Map
Figure 14: Visualization of result in OMA Val set.
Figure 15: Visualization of result in OMA Test set.
H.2 Result Compare
Fig.14 illustrates comparisons of model ground truth on KNN, HMM, MAT and MAT w/ postprocess. The top rows of (d) and (e) exhibit our post-processing module’s effectiveness. MAT predictions initially display incorrect topological connections where roads are mistakenly linked (marked with red circles). Our postprocessing, which utilizes topology-aware beam searching, rectifies this by eliminating non-sequential transitions and reconstructing precise topological paths. The second row demonstrates MAT’s superior handling of complex topologies. Although HMM targets sequential path associations, its single path paradigm often underperforms in complex topologies with intersections. In contrast, our model uses spatial attention to grasp global information and cross-path associations, facilitating adaptive learning of complex topological patterns for accurate connectivity inference. The third row showcases our model’s improved ability to localize associations. Using path-aware attention, the model emphasizes detailed extraction of local features along paths. This targeted local perception ensures precise associations at challenging points, such as junctions, where HMM is typically short due to limited contextual understanding.
Fig. 15 illustrates a comparison of ground truth results for KNN, HMM, MAT, and MAT with post-processing. Significantly, the visualization demonstrates that our model excels in map association in noisy scenarios with inaccurate centerline predictions, surpassing KNN and HMM by integrating the complementary benefits of global association (SA) and local detail refinement (PA).
Figure 16: Visualization of failed cases.
H.3 Failed Cases
H.4 Generalization on Different OP Generators
To further validate the robustness of MAT against varying noise patterns inherent to different map generation paradigms, we visualize the association results on the OMA Test set using three distinct baselines: MapTR (liao2023maptr), MapTRv2 (maptrv2), and SeqGrowGraph (SSG) (xie2025seqgrowgraph).
As illustrated in Figure 17, each generator exhibits unique geometric and topological characteristics:
  • Row 1 (MapTR): Often produces fragmented centerlines with localized geometric jitter, creating a challenge to ensure topological continuity.
  • Row 2 (MapTRv2): Offers improved geometric stability, but still suffers from occasional detection gaps and instance discontinuities.
  • Row 3 (SeqGrowGraph): Generates highly connected graphs via expansion, which reduces fragmentation, but may introduce erroneous topological links (over-connection) in complex intersections.
Despite these substantial domain gaps, MAT consistently establishes accurate associations (indicated by the consistent coloring of lane instances that match the SD map topology) across all three inputs. This qualitative evidence reinforces the quantitative results in Tab. 1, demonstrating that our proposed Path-Aware and Spatial Attention mechanisms effectively generalize across disparate upstream perception generators without requiring generator-specific fine-tuning.
Figure 17: Visualization of MAT association results across different Online Perception (OP) map generators on the OMA Test set. Rows correspond to (1) MapTR, (2) MapTRv2, and (3) SeqGrowGraph (SSG). The columns display diverse scenarios including intersections and straight roads. Consistent coloring between lane segments indicates that MAT successfully associates fragmented or noisy OP elements with the correct semantic road instances from the SD map, demonstrating strong cross-generator robustness.
Appendix I Implement Details
I.1 Model Settings
@@
Tab. 22 summarizes the architectural configurations of our proposed MAT variants (MAT-T and MAT-L). All variants adopt identical channel dimensions , attention head counts, spatial curve orders and hybrid attention mechanisms combining spatial attention (SA) and path-aware attention (PA). In particular, parameters such as patch sizes, MLP ratios (matching spatial curve orders), and stochastic depth rates are uniformly inherited across architectures, reflecting ablation study results that optimized these values for balanced accuracy-latency trade-offs. A distinctive design choice lies in the shuffling strategy, where MAT-T/L progressively refine the shuffle order to enhance token mixing in spatial attention, aligning with their increasing computational budgets. This structured configuration hierarchy enables systematic evaluation of model capacity versus efficiency, as validated by the ascending La / ms metrics (34 $\rightarrow$ 70) corresponding to deeper transformer layers.
Table 22: Model settings.
Parameter MAT-T MAT-L
Blocks [2, 2, 2, 2, 2] [4, 4, 4, 12, 4]
@@
Latency/ms 34 70
Table 23: Train Configuration and Data augmentations.
@@
learning rate 1e-4 epochs 50
criteria CrossEntropy, CTC Loss warmup epochs 2
@@
I.2 Train Configuration
The details of the implementation are summarized in Tab. 23. Our training protocol uses the AdamW optimizer with a base learning rate of $1\times 10^{-4}$ and a cosine learning rate decay, operating in batch size of 128. The weight decay regularization is set to $5\times 10^{-3}$. The training process spans 50 epochs with a 2-epoch warm-up phase for learning rate initialization. Model optimization combines CrossEntropy loss for classification tasks and CTC loss for sequence alignment objectives.
I.3 Data Augmentations
For data augmentation as shown in Tab. 23, we implement a series of randomized transformations that include axis-aligned rotation around the z-axis with $\pm 1^{\circ}$ angular variation at a 50% application probability, isotropic scaling within the range $[0.9,1.1]$, random flipping with equal probability 50%, point cloud jittering characterized by $\sigma=0.005$ and a clip limit of 0.02, and grid sampling with spatial discretization parameters set to $0.1,0.1,\pi/16]$. These enhancement strategies were systematically validated through ablation studies to optimize the balance between model accuracy and computational efficiency while ensuring robustness to input variations.